OpenStudy (anonymous):

2^(2x)-3*(2^x)=-2 solve for x

OpenStudy (anonymous):

Assume 2^x=a ............(1) a^2-3a=-2 or a^2-3a+2=0. ..............(2) This is quadratic equation, solve for a. Once you find a, you can find x from (1).

OpenStudy (anonymous):

So to begin solving the quadratic equation, would I divide both sides by three, so as to isolate the variables to the left?

OpenStudy (anonymous):

Oh, nevermind- I'd get one side to equal zero, as you showed, then work from there.

OpenStudy (anonymous):

I think it is an easy to factor quadratic equation. Just factor it.

OpenStudy (anonymous):

(a-3)(a+1)... is this correct?

OpenStudy (anonymous):

Nope, try again :)

OpenStudy (anonymous):

a(a-3)?

OpenStudy (anonymous):

No, your previous answer almost correct. Here's how to think it: From the quadratic eq, a=1, b=-3, c=2 find two number X and Y, where X*Y = a*c=2 and X+Y=b=-3 Give up?

OpenStudy (anonymous):

(a-2)(a-1)

OpenStudy (anonymous):

Yup, finally. Just do the rest.

OpenStudy (anonymous):

(2x-2)(2x-1) ?

OpenStudy (anonymous):

No, (a-2)(a-1)=0 means a=2 or a=1 (we have two answer here) 2^x=2 or 2^x=1 x=2log2 or x=0

OpenStudy (anonymous):

okay! Great! One more question: In the beginning, when we substituted 2^x for a, why did the first 2^x become a^2, while the one which was multiplied by the 3 was just "a"? Was this do to the quadratic form?

OpenStudy (anonymous):

Thank you sooo much by the way!!! :)

OpenStudy (anonymous):

2^(2x)=(2^x)^2=(2^2)^x This is just exponential rule. Try it with numbers, you'll understand.